# Systematic Uncertainties in the Measurement of Neutron lifetime Using Lunar Prospector Neutron Spectrometer

### v2

@@ Abstract
The lifetime of free neutrons measured in the lab has a long standing disparity of $\sim$9 s. A space-based technique has recently been proposed to independently measure the neutron lifetime using interactions between the galactic cosmic rays and a low atmosphere planetary body. This technique has not produced competitive results yet due to constraints of non-optimized data that contain large systematic errors. We use data from the neutron spectrometer on-board NASA’s Lunar Prospector, and quantify the effects of lunar sub-surface temperature and the composition on the measurement of neutron lifetime. We use the HeCd and HeSn neutron spectrometer data when the spacecraft was in a highly elliptical orbit during the orbit insertion period. We report the neutron lifetime using lunar composition maps at different resolutions: 5 Prettyman et al. (2006) and 2 re-binned Wilson et al. (2021) maps to be 777.59$\pm$11.71 s and 739.64$\pm$10.83 s respectively. Further, for the 20 map, we apply both a constant equatorial and a latitude-dependent temperature model and report 738.63$\pm$10.75 s and 767.33$\pm$11.17 s respectively. We report the weighted average of the four measurements to be $\tau_{n}=754.72^{+34.58}_{-25.91}$ s.
@@ I Introduction
There are primarily two different methods of measuring neutron lifetime: beam and bottle experiments. The beam experiment measures the appearance of protons from the decay of a beam of neutrons (Wietfeldt and Greene, 2011). Robson (1951) made one of the first measurements of the neutron lifetime using the beam method, with the most precise recent measurement being $\tau_{n}=886.3\pm 3.2~{}s$ by Nico et al. (2005). Further, a more precise value is expected with $<$1 s of uncertainty in the next few years (Hoogerheide et al., 2019). The bottle method measures the disappearance of neutrons in a confined system as a function of time. Kosvintsev et al. (1980) gave the proof of principle in 1980 and reported $\tau_{n}=875\pm 95~{}s$. The uncertainties in the measurement were significantly improved within a decade to $\tau_{n}=887.6\pm 3~{}s$ as reported by Mampe et al. (1989) and has been getting better with the most recent measurement being $\tau_{n}=877.75\pm 0.28_{\text{stat}}+0.22/{-0.16_{\text{syst}}}~{}\textsl{s}$ by Gonzalez et al. (2021). Although we have these measurements upto an accuracy of 1%, they differ by over $5\sigma$. One or both measurement techniques must have complex, systematic uncertainties that have yet to be resolved.
@@
A third method of measuring neutron lifetime is now actively being explored. A space-based measurement technique makes use of the free neutrons that are naturally produced by interactions between Galactic Cosmic Rays (GCRs) and a planetary body, such as the Moon. The lowest energy ($<$ 0.5 eV) of these neutrons have a velocity less than 10 km/s, and can thus undergo beta decay in the distances reasonably sampled by an orbiting satellite. The neutron lifetime can then be measured based on the disappearance of these neutrons with distance from the planetary body. These same neutrons have been measured by many planetary science missions because they provide insight into the surface composition of planetary bodies. Thus while no dedicated space-based neutron lifetime measurement has been performed, initial studies are possible with existing planetary science data. A feasibility study of this method was first done by Wilson et al. (2020), estimating the $\tau_{n}=780\pm 70~{}s$ using flybys of Venus and Mercury by NASA’s MESSENGER spacecraft. A more recent measurement of $\tau_{n}=887\pm 14~{}s$ has been reported by Wilson et al. (2021) using data from the Lunar Prospector (LP) mission.
In this paper, we consider uncertainties in the neutron lifetime measurement due to uncertainties in the neutrons emitted from the lunar surface. The energy spectrum of the neutrons that leak from a planetary surface is dependent on the local surface composition (Mesick et al., 2018), which is why neutron measurements are an effective tool in planetary science. Specifically, we study the effects of temperature of the lunar subsurface and the resolution of composition maps on the computation of the neutron lifetime. We consider surface compositions derived from orbital measurements at three spatial resolutions. At the smallest spatial resolution (2 in latitude and longitude) we limit ourselves to considering five distinct compositions, following the work of Wilson et al. (2021). At coarser resolutions (5 and 20 latitude with roughly equal spatial size in longitude) we use compositions derived from LP Gamma-ray Spectrometer (GRS) measurements, which are sampled at similar spatial and depth scales as the LP Neutron Spectrometer (LPNS). As the thermal neutron emission may be impacted by the temperature of the lunar subsurface, we consider both constant temperature and a latitude-dependent temperature model specifically developed for the analysis of LPNS data. Starting from these lunar surface definitions, we construct a simulation chain that models LPNS data as a function of neutron lifetime during the LP early-mission high-altitude orbits. We compare the simulated neutron current to the measured LPNS data to arrive at the neutron lifetime. We compare the values derived from different surface treatments and discuss the uncertainty in the measurement.
@@
We note that the LP mission was a planetary science mission and was not designed for a neutron lifetime measurement. Thus the uncertainty reported here using a very limited non-optimized data set should not be expected to compare with the accuracy and precision of current beam or bottle measurement techniques, but instead, this analysis serves as a study of systematic uncertainties for the space-based technique for future neutron lifetime measurements.
@@ II.1 Neutron Detector
LP consists of two geometrically identical neutron detector proportional counter tubes, each of 20 cm in length and 5.7 cm in diameter (Feldman et al., 2004). The tubes are filled with 3He gas at 10 atm pressure and covered with a 0.63 mm thick sheet of either Cd or Sn. The Cd absorbs low energy neutrons, so the “HeCd” detector is sensitive to neutrons of energy $>0.4$ eV (epithermal neutrons) while the “HeSn” detector is sensitive to neutrons of all energies (thermal and epithermal neutrons). The data product from these detectors is a 32-bin pulse amplitude histogram that is produced at a 32 second cadence. We sum the bins in the second half of the pulse height spectra, bins corresponding to higher energy deposition events in the detector to reduce the contribution of non-neutron background interactions, corresponding to energy bins 17 to 32. This produces the neutron counts per 32 seconds data product for the HeCd and HeSn detectors and the corresponding uncertainties in measurements. This choice was made to avoid the background noise and to achieve higher signal-to-noise in the pulse height spectrum (Maurice et al., 2004). To get the thermal neutron counts (energy range of $<$0.4 eV), we subtract the HeCd count rates from the HeSn count rates.
@@ II.2 Simulating Surface Composition & Temperature
Wilson et al. (2021) derived five lunar compositions from Peplowski et al. (2016) and Jolliff et al. (2000), which are re-binned into a 2 resolution map for this work. These compostions are the procellarum KREEP terrane (PKT), South Pole Aitken terrane (SPA), feldspathic highland terrane (FHT), pure anorthosite terrane (PAN) and non-PKT maria (nPKT). For elemental mass fraction of these composition, refer to Table II in Wilson et al. (2021). Although this map has been used to estimate the value of neutron lifetime, it is a simplified map, and fails to capture the complexities of the lunar terrain. Here, in addition to the five listed compositions, we also model the neutron current from the lunar surface at two more resolutions derived from Prettyman et al. (2006) - (i) a 5 resolution map showing 1790 different compositions and (ii) a 20 resolution map showing 114 different compositions. The 5 binned composition map comes from the NASA Planetary Data System (PDS) (LPP, 2002) and contains weight percent (wt%) or ppm of Mg, O, Al, Si, Ca, Ti, Fe, K, Th, and U. Additional higher resolution (2) binned maps for H and Sm are included from the PDS. To form a complete set, the H and Sm were resampled to 5 and 20 resolutions. Further, Gd, an important neutron absorbing element, was included by scaling the amount of Sm by a constant factor of 1.3, adopted from Hidaka et al. (2000).
@@ II.3 Neutron Emission Simulations
To simulate the GCRs we used the particle flux equations in Usoskin et al. (2017) and the reported mean solar modulation during the orbital insertion of LP (Jan 1998) of 505 MV, which results in a GCR proton flux of 3.575 particles/cm2-s and $\alpha$ particle flux of 0.323 particles/cm2-s. We simulate only protons, but account for the contribution of $\alpha$ flux using the scaling from McKinney et al. (2006) of a factor of 3.8 more neutrons produced per incident $\alpha$ than per incident proton. Thus the total equivalent omni-directional proton flux is $3.575+3.8\times 0.323=4.8$ p/cm2-s. Since our MCNP simulations start with a $2\pi$ surface instead of total omni-directional flux, the incident flux is a factor of 4 lower (McKinney et al., 2006, Mesick et al., 2018), resulting in an incident simulated proton flux of 1.2 particles/cm2-s.
We tally the neutrons emitted after the interaction of GCRs with the lunar surface using MCNP current tallys and normalize for source flux (1.2 particles/cm2-s) and tally area. We also tested the angular distributions of a few compositions and found the neutron current to be independent of the cos($\theta$)1.25 relationship, $\theta$ being angle of neutron travel relative to the tally surface. This is unlike the flux tally which needs additional appropriate solid angle normalization. To keep the simulations computationally minimal, we do not use the angle-dependent outputs. For each of the compositions tested, the normalized neutron current integrated over the thermal peak ($<$0.4 eV, when neutrons have slowed to an energy/velocity equal to a temperature dependent thermal equilibrium) is noted to be $\approx$0.2 particles/cm2-s.
@@
MCNP is designed to compute neutron current by creating surfaces through which neutrons cross after they are released from a planetary surface, such as the Moon. This approach assumes that the neutron escapes the gravitational field of the planetary body. However, low-energy neutrons are gravitationally-bound and can return, interact with the surface, and be re-emitted, which increases the neutron current at low energies(Feldman et al., 1989). It may therefore be necessary to account for the low-energy neutrons with an energy less than 0.0295 eV, that do not escape lunar gravity. These effects can be accounted for using the FIELD card in MCNP6.2+ and MCNPX, which were implemented based on particle physics discussed in Feldman et al. (1989). However, we tested the effects of gravitationally bound downward travelling neutrons using the 5 map and found that MCNP 6.2 uses the incorrect value of $\tau_{n}$ resulting in an over-prediction of the gravity effects. We estimate the gravitational effect to be less than 0.5% on the total thermal neutron flux for lunar compositions with the low water content (tens of ppm) used in this study. We therefore disabled the FIELD card in our simulations.
@@ II.4 Propagation Model
Lunar Prospector was spin-stabilized, rotating around its axis with a period of roughly 5 seconds, which is much shorter than the 32 second integration time of the primary data set. The detector response was thus computed spin-averaged around the spacecraft rotation axis. The spacecraft bus itself shadows the detectors in some of the geometries. This effect was corrected empirically in the LPNS primary mission data (Maurice et al., 2004) rather than being modeled. However, those results are not applicable to the orbital insertion data used in this study. Thus, this effect is not corrected in our analysis. The results of Maurice et al. (2004) suggest that the impact of this effect is $\approx$ 30% shadowing for HeCd at the equator and reduces steadily as the spacecraft moves towards the poles. HeSn is expected to experience similar effects with an 180 offset. The final spin-averaged detector responses are shown in Figure 3 for each detector, expressed as the detector effective area, i.e., the size of an equivalent perfectly efficient detector. Neutron flux was convolved with the detector response to obtain the detector count rates. The simulated detector count rate, for each of the two detectors, is thus the product of the incident neutron flux and effective area integrated over all energies and angles.
@@ III.2 Effects of surface temperature
where $f_{C}$ and $f_{V}$ represent normalized neutron current for constant equatorial temperature and latitude dependent subsurface temperature MCNP outputs respectively, and $\sigma_{C}$ and $\sigma_{V}$ represent corresponding standard deviations.
Figure 7 shows the difference in integrated thermal neutrons, correlation coefficients for neutrons of energy $<0.4~{}eV$, and the corresponding residuals in thermal neutrons for each of the two temperature conditions. For the calculation of correlation coefficients, a nominal value of $\tau_{n}$ = 880 s is assumed. Table 2 gives the RMS in the residuals shown in Figure 7. We note that, although a simple case of constant lunar equatorial temperature results in lower RMS, a latitude dependent temperature condition is realistic and such models are therefore necessary to capture the effects of sub-surface lunar temperatures in the measurement of neutron lifetime.
Table 2: Residuals in thermal neutrons between LP data and models obtained using different lunar subsurface temperatures, evaluated for the 20 map.
@@
Figure 7: (Top left:) Difference in integrated thermal neutrons ($E<0.4$ MeV) for two different lunar sub-surface temperature conditions - (i) constant equatorial temperature and (ii) latitude dependent temperature, computed for the 20 resolution map. Higher neutron current is seen in at the poles when temperature is set same as that of the equator. (Top right:) Correlation coefficients between the two cases. Highest correlation and neutron currents are seen at equators and lowest correlation and neutron currents at the poles where it is colder, indicating the necessity of accounting for the varying surface temperatures in the propagator model.

### v3

@@ Abstract
The lifetime of free neutrons measured in the lab has a long standing disparity of $\sim$9 s. A space-based technique has recently been proposed to independently measure the neutron lifetime using interactions between the galactic cosmic rays and a low atmosphere planetary body. This technique has not produced competitive results yet due to constraints of non-optimized data that contain large systematic errors. We use data from the neutron spectrometer on-board NASA’s Lunar Prospector, and study two large systematics in the measurement of neutron lifetime: the lunar sub-surface temperature and the lunar surface composition. We use the HeCd and HeSn neutron spectrometer data when the spacecraft was in a highly elliptical orbit during the orbit insertion period. We report the neutron lifetime using four different models that each have different choices of surface temperature and composition. 5 Prettyman et al. (2006) and 2 re-binned Wilson et al. (2021) maps result in 777.6$\pm$11.7 s and 739.6$\pm$10.8 s respectively. For the 20 map, constant equatorial and a latitude-dependent temperature model result in 738.6$\pm$10.8 s and 767.3$\pm$11.2 s respectively. Increasing the complexities of the models accounting for the systematic effects increase the measured lifetime. However, the reported measurements are not competitive with the laboratory results due to large unaccounted systematics resulting from non-optimized measurements and modeling assumptions. This work serves as a study of systematic uncertainties for future neutron lifetime measurements using the space-based technique. We estimate the effect on the lifetime from the choice of temperature model to be to be 28.7 $\pm$ 15.5 s, and choice of compositional map (for 20 and 5 maps) to be 10.3 $\pm$ 12.2 s.
@@ I Introduction
There are primarily two different methods of measuring neutron lifetime: beam and bottle experiments. The beam experiment measures the appearance of protons from the decay of a beam of neutrons (Wietfeldt and Greene, 2011). Robson (1951) made one of the first measurements of the neutron lifetime using the beam method, with the most precise recent measurement being $\tau_{n}=887.7\pm 1.2_{stat}\pm 1.9_{syst}~{}s$ by Yue et al. (2013). Further, a more precise value is expected with $<$1 s of uncertainty in the next few years (Hoogerheide et al., 2019). The bottle method measures the disappearance of neutrons in a confined system as a function of time. Kosvintsev et al. (1980) gave the proof of principle in 1980 and reported $\tau_{n}=875\pm 95~{}s$. The uncertainties in the measurement were significantly improved within a decade to $\tau_{n}=887.6\pm 3~{}s$ as reported by Mampe et al. (1989) and has been getting better with the most recent measurement being $\tau_{n}=877.75\pm 0.28_{\text{stat}}+0.22/{-0.16_{\text{syst}}}~{}\textsl{s}$ by Gonzalez et al. (2021). The most precise beam and bottle methods differ by over $4\sigma$ rajan2020meta. This discrepancy is most likely attributed to potentially unresolved systematic uncertainties in one or both measurement techniques.
@@
A third method of measuring neutron lifetime is now actively being explored. A space-based measurement technique makes use of the free neutrons that are naturally produced by interactions between Galactic Cosmic Rays (GCRs) and a planetary body, such as the Moon. The lowest energy ($<$ 0.5 eV) of these neutrons have a velocity less than 10 km/s, and can thus undergo beta decay in the distances reasonably sampled by an orbiting satellite. The neutron lifetime can then be measured based on the disappearance of these neutrons with distance from the planetary body. These same neutrons have been measured by many planetary science missions because they provide insight into the surface composition of planetary bodies. Thus while no dedicated space-based neutron lifetime measurement has been performed, initial studies are possible with existing planetary science data. A feasibility study of this method was first done by Wilson et al. (2020), estimating the $\tau_{n}=780\pm 70~{}s$ using flybys of Venus and Mercury by NASA’s MESSENGER spacecraft. A more recent measurement of $\tau_{n}=887\pm 14~{}s$ has been reported by Wilson et al. (2021) using data from the Lunar Prospector (LP) mission. It is pointed out in Wilson et al. (2021) that the biggest sources of systematic uncertainties in such measurements are expected to be the choice of models describing the lunar composition and latitudinal temperature variations that dictate the thermal neutron flux.
@@
The LP mission was a planetary science mission and was not designed for a neutron lifetime measurement; therefore, we do not expect the uncertainty reported here to compare with the accuracy and precision of current beam or bottle measurement techniques. This analysis serves as a study of systematic uncertainties for the space-based technique for future neutron lifetime measurements. While several sources of systematic uncertainties exist in the space-based measurement technique, in this paper, we consider uncertainties in the neutron lifetime measurement due to the choice of a surface composition model and a temperature model. We test four different models that capture these systematics and report a neutron lifetime for each case. The energy spectrum of the neutrons that leak from a planetary surface is dependent on the local surface composition (Mesick et al., 2018), which is why neutron measurements are an effective tool in planetary science.
Specifically, we study the effects of temperature of the lunar subsurface and the resolution of composition maps on the computation of the neutron lifetime. We consider surface compositions derived from orbital measurements at three spatial resolutions. At the smallest spatial resolution (2 in latitude and longitude) we limit ourselves to considering five distinct compositions, following the work of Wilson et al. (2021). At coarser resolutions (5 and 20 latitude with roughly equal spatial size in longitude) we use compositions derived from LP Gamma-ray Spectrometer (GRS) measurements, which are sampled at similar spatial and depth scales as the LP Neutron Spectrometer (LPNS). As the thermal neutron emission may be impacted by the temperature of the lunar subsurface, we consider both constant temperature and a latitude-dependent temperature model specifically developed for the analysis of LPNS data. Starting from these lunar surface definitions, we construct a simulation chain that models LPNS data as a function of neutron lifetime during the LP early-mission high-altitude orbits. We compare the simulated neutron current to the measured LPNS data to arrive at the neutron lifetime. We compare the values derived from different surface treatments and discuss the uncertainty in the measurement.
@@ II.1 Neutron Detector
LP consists of two geometrically identical neutron detector proportional counter tubes, each of 20 cm in length and 5.7 cm in diameter (Feldman et al., 2004). The tubes are filled with 3He gas at 10 atm pressure and covered with a 0.63 mm thick sheet of either Cd or Sn. The Cd absorbs low energy neutrons, so the “HeCd” detector is sensitive to neutrons of energy $>0.4$ eV (epithermal neutrons) while the “HeSn” detector is sensitive to neutrons of all energies (thermal and epithermal neutrons). The data product from these detectors is a 32-bin pulse amplitude histogram that is produced at a 32 second cadence. We sum the bins in the second half of the pulse height spectra, bins corresponding to higher energy deposition events in the detector to reduce the contribution of non-neutron background interactions, corresponding to energy bins 17 to 32. This produces the neutron counts per 32 seconds data product for the HeCd and HeSn detectors and the corresponding uncertainties in measurements. This choice was made to avoid the background noise and to achieve higher signal-to-noise in the pulse height spectrum (Maurice et al., 2004). The background rates in this data were assessed as the average measured rate far from the Moon in the days before lunar orbital insertion. These background rates are 134.9 and 127.8 counts/32-s for the HeCd and HeSn detectors, respectively. Backgrounds are the result of GCR spallation on the spacecraft itself, and so they decrease at lower altitudes as the Moon blocks incident GCRs. The decrease in background is linear with the solid angle subtended by the Moon as a function of altitude, which is 35% at the lowest altitudes in this dataset. Background-subtracted rates are shown in the top panel of Figure 4. To get the thermal neutron counts (energy range of $<$0.4 eV), we subtract the HeCd count rates from the HeSn count rates.
@@ II.2 Simulating Surface Composition & Temperature
Wilson et al. (2021) derived five lunar compositions from Peplowski et al. (2016) and Jolliff et al. (2000), which are re-binned into a 2 resolution map for this work. These compostions are the procellarum KREEP terrane (PKT), South Pole Aitken terrane (SPA), feldspathic highland terrane (FHT), pure anorthosite terrane (PAN) and non-PKT maria (nPKT). For elemental mass fraction of these composition, refer to Table II in Wilson et al. (2021). Although this map has been used to estimate the value of neutron lifetime, it is a simplified map, and fails to capture the complexities of the lunar terrain. Here, in addition to the five listed compositions, we also model the neutron current from the lunar surface at two more resolutions derived from Prettyman et al. (2006) - (i) a 5 resolution map showing 1790 different compositions and (ii) a 20 resolution map showing 114 different compositions. The 5 binned composition map comes from the NASA Planetary Data System (PDS) (LPP, 2002) and contains weight percent (wt%) or ppm of Mg, O, Al, Si, Ca, Ti, Fe, K, Th, and U. Additional higher resolution (2) binned maps for H and Sm are included from the PDS. To form a complete set, the H and Sm were resampled to 5 and 20 resolutions. Further, Gd, an important neutron absorbing element, was included by scaling the amount of Sm by a constant factor of 1.3, adopted from Hidaka et al. (2000). We note that, although both the 5 and 20 maps are obtained from the PDS, the 20 map is not directly derivable from the 5 map. Since the post-processing involves various steps including spatial averaging, interpolation, noise reduction among others, we use the available published maps.
@@ II.3 Neutron Emission Simulations
To simulate the GCRs we use the particle flux equations in Usoskin et al. (2017) and the reported mean solar modulation during the orbital insertion of LP (Jan 1998) of 505 MV, which results in a GCR proton flux of $J_{p}$=3.575 particles/cm2-s and $\alpha$ particle flux of $J_{alpha}$=0.323 particles/cm2-s. We simulate only protons, but account for the contribution of $\alpha$ flux using the scaling from McKinney et al. (2006) of a factor of $R=3.8$ more neutrons produced per incident $\alpha$ than per incident proton. Thus the total equivalent omni-directional proton flux $J_{eq,p}=J_{p}+RJ_{alpha}$ is 4.8 p/cm2-s. Since our MCNP simulations start with a $2\pi$ surface instead of total omni-directional flux, the incident flux is a factor of 4 lower (McKinney et al., 2006, Mesick et al., 2018), resulting in an incident simulated proton flux of $J_{p,inc}$= 1.2 particles/cm2-s.
We tally the neutrons emitted after the interaction of GCRs with the lunar surface using MCNP current tallys and normalize for source flux ($J_{p,inc}$) and tally area. We tested the angular distributions of a few widely-varying compositions and found that the angular distributions did not vary with composition. Most MCNP simulations were then performed without tallying the emission as a function of angle, which improved the convergence of the simulations with lower computational requirements. For each of the compositions tested, the normalized neutron current integrated over the thermal peak ($<$0.4 eV, when neutrons have slowed to an energy/velocity equal to a temperature dependent thermal equilibrium) is noted to be $\approx$0.2 particles/cm2-s.
@@
MCNP is designed to compute neutron current by creating surfaces through which neutrons cross after they are released from a planetary surface, such as the Moon. This approach assumes that the neutron escapes the gravitational field of the planetary body. However, low-energy neutrons are gravitationally-bound and can return, interact with the surface, and be re-emitted, which increases the neutron current at low energies(Feldman et al., 1989). It may therefore be necessary to account for the low-energy neutrons with an energy less than 0.0295 eV, that do not escape lunar gravity. These effects can be accounted for using the FIELD card in MCNP6.2+ and MCNPX, which were implemented based on particle physics discussed in Feldman et al. (1989). However, we tested the effects of gravitationally bound downward traveling neutrons using the 5 map and found that MCNP 6.2 does not decay neutrons, resulting in an over-prediction of the gravity effects. We estimate the effect of ignoring gravitational bound neutrons to be less than 0.5% on the total thermal neutron flux for lunar compositions with the low water content (tens of ppm) used in this study. We therefore disabled the FIELD card in our simulations.
@@ II.4 Propagation Model
Lunar Prospector was spin-stabilized, rotating around its axis with a period of roughly 5 seconds, which is much shorter than the 32 second integration time of the primary data set. The detector response was thus computed spin-averaged around the spacecraft rotation axis. The spacecraft bus itself shadows the detectors in some of the geometries. This effect was corrected empirically in the LPNS primary mission data (Maurice et al., 2004) rather than being modeled. However, those results are not applicable to the orbital insertion data used in this study. Thus, this effect is not corrected in our analysis. The results of Maurice et al. (2004) suggest that the impact of this effect is between 0% and $\approx$ 30% shadowing depending on the orientation of the spacecraft relative to the lunar surface. The final spin-averaged detector responses are shown in Figure 3 for each detector, expressed as the detector effective area, i.e., the size of an equivalent perfectly efficient detector. Neutron flux was convolved with the detector response to obtain the detector count rates. The simulated detector count rate, for each of the two detectors, is thus the product of the incident neutron flux and effective area integrated over all energies and angles.
@@ III.2 Effects of surface temperature
where $f_{C}$ and $f_{V}$ represent normalized neutron current for constant equatorial temperature and latitude dependent subsurface temperature MCNP outputs respectively, and $\sigma_{C}$ and $\sigma_{V}$ represent corresponding standard deviations; and $\mu_{C}$ and $\mu_{V}$ are mean of $f_{C}$ and $f_{V}$ respectively.
Figure 7 shows the difference in integrated thermal neutrons, correlation coefficients for neutrons of energy $<0.4~{}eV$, and the corresponding residuals in thermal neutrons for each of the two temperature conditions. For the calculation of correlation coefficients, a nominal value of $\tau_{n}$ = 880 s is assumed. Table 2 gives the RMS in the residuals shown in Figure 7. We note that the simple case of constant lunar equatorial temperature results in lower RMS, showing better agreement with LP data compared to latitude dependent temperature model. However, we note that the latitude dependent model is a realistic representation of the lunar subsurface temperatures and such models are therefore necessary in the measurement of neutron lifetime.
Table 2: Residuals in thermal neutrons between LP data and models obtained using different lunar subsurface temperatures, evaluated for the 20 map. A lower RMS here indicates better agreement to LP data.
@@
Figure 7: (Top left:) Difference in integrated thermal neutrons ($E<0.4$ MeV) for two different lunar sub-surface temperature conditions - (i) constant equatorial temperature (latitude independent) and (ii) latitude dependent temperature, computed for the 20 resolution map. Higher neutron current is seen in at the poles when temperature is set same as that of the equator. (Top right:) Correlation coefficients between the two cases. Highest correlation and neutron currents are seen at equators and lowest correlation and neutron currents at the poles where it is colder, indicating the necessity of accounting for the varying surface temperatures in the propagator model. (Bottom left:) Residuals in thermal neutrons between LP data and models for latitude independent temperature model and (Bottom right:) residuals in thermal neutrons between LP data and models for latitude dependent temperature model. For both the cases, the data at higher altitude sees a lower level of residuals than at lower altitudes.